# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### v2

@@ G.3 Results when using $R=2$ for all algorithms
Figure 36: Bernoulli GLM model with $R=2$ for all algorithms. Posterior plots obtained with SeMPLE and obtained after 5,000 and 10,000 model simulations.
@@ G.4 Results when using $R=10$ for SNL/SNPE-C and $R=2$ for SeMPLE
We executed the experiment anew, but this time by allocating the $10^{4}$ model simulations uniformly across 2 rounds for SeMPLE and 10 rounds for SNL and SNPE-C (as per SBIBM defaults). Results are in Figure 39 and Figure 40. Once more, we recommend the reader to consider the Wasserstein distances and not C2ST for this example, since that posterior is unimodal, as motivated in the previous section. The distances show that on average SeMPLE is as accurate as SNL but not before $10^{4}$ model simulations are executed, while SNPE-C shows a large variability in the results. Once more, SeMPLE comes with the benefit of a much faster inference (Figure 40: in this case SeMPLE is 26 times faster than SNPE-C and 74 times faster than SNL.
@@ Appendix H Ornstein-Uhlenbeck
This makes the Ornstein-Uhlenbeck model one of the very few examples of SDEs where it is trivial to evaluate the exact likelihood function. The likelihood function for the parameters $\bm{\theta}=(\alpha,\beta,\sigma)$ is written as (due to the Markovanity of the solution of an SDE)
@@ H.3 Results when using $R=4$ for all algorithms
In terms of runtime, Figure 44 shows that SeMPLE has a much much smaller runtime compared to both SNL and SNPE-C, even if were to include the 16 minutes of BIC computations into account. Especially SNPE-C has much longer runtimes with this model, combined with the worst performance metric results out of all three algorithms. Plots of the posterior marginals and posterior draws from the last round (r=4) of each algorithm are in Figures 45-46, for each of the 10 independent runs. Acceptance rates of the Metropolis-Hastings step of SeMPLE are in Figure 47, showing that, particularly for the third and fourth round, the acceptance rate is fairly high (compatibly with being an MCMC algorithm).
@@ I.2 Inference setup
Prior to running the inference, for each of the nine summary statistics we computed trimmed means, where $1.25\%$ of the values were trimmed from each end and corresponding trimmed standard deviations, obtained from 10,000 prior predictive simulations, and we used these to standardise both the observed summaries and the summaries simulated during the inference. For each SeMPLE run we set a simulation budget of 30 0003000030\,00030 000 uniformly distributed across 3 algorithm rounds. The number of mixture components is set to $K=15$ (a number which was picked without any specific tuning) and we set the covariance matrices in GLLiM to be full unconstrained matrices, all different across the $K$ components. Moreover, we do not inflate the covariance in the Metropolis-Hastings step (ie we set $\gamma=1$). The mixture probability threshold to remove negligible mixture components during the SeMPLE run was set to $0.005$.
For a MJP Lotka-Volterra model, we did not find a ready implementation in SBIBM (there, an ODE-based Lotka-Volterra model is provided, not a MJP version), and this is likely because with such a model it is impossible to perfectly control the computational effort, as for a given value of $\theta$ the time spent to complete a trajectory is random. For this reason, we found more practical to compare SeMPLE with approximate Bayesian computation (ABC) experiments using the code from Picchini & Tamborrino (2022). For SMC-ABC, we used the same observed data and the same standardised summary statistics as with SeMPLE. We consider sequential Monte Carlo ABC (SMC-ABC), and also use the posterior from SMC-ABC as “reference posterior”. SMC-ABC is the state-of-art approach to ABC inference and is the algorithm of choice in recent inference platforms such as ABCpy (Dutta et al., 2017) and pyABC (Schälte et al., 2022). We consider two versions of SMC-ABC that differ in terms of proposal function: the first one is the “classic” SMC-ABC, in that it uses as proposal distribution a Gaussian random walk of the type $\theta^{**}\sim N(\theta^{*},2\Sigma)$ (Beaumont et al., 2009, Filippi et al., 2013), where $\theta^{*}$ is a parameter “particle” randomly sampled from the most recent set of “parameter particles” (these providing an ABC approximation to the posterior), and $\Sigma$ is the empirical weighted covariance matrix of these “parameter particles”. This first sampler is named standard in Picchini & Tamborrino (2022), as it is the most common SMC-ABC proposal sampler, as found in most literature and in available software packages. A second SMC-ABC sampler is the one denoted blockedopt in Picchini & Tamborrino (2022), which is also a Gaussian proposal sampler but not of random-walk type, and is much more resource-effective, as it more rapidly “guide” the parameters towards the bulk of the posterior region (see Picchini & Tamborrino, 2022 for details). We ran both SMC-ABC versions with 1,000 “particles”.
@@ I.3 Results
We now consider a model with 20 parameters and a challenging posterior characterised by a strong correlation between some of the parameters. This case study is particularly interesting, as the likelihood only provides location information about the unknown parameter while the dependence structure in the posterior comes mostly from the prior. For this reason, the posterior dependence changes direction depending on whether the likelihood locates the posterior in the left or right tail of the prior (see Nott et al., 2018 for a graphical illustration). This case study was analysed in an ABC context in Li et al. (2017) and Picchini & Tamborrino (2022). The model assumes $\bm{y}=(y_{1},...,y_{d_{\theta}})$ drawn from a $d_{\theta}$-dimensional Gaussian $\bm{y}\sim\mathcal{N}(\bm{\theta},\bm{\Psi})$, with $\bm{\theta}=(\theta_{1},...,\theta_{d_{\theta}})$ and diagonal covariance matrix $\Psi=\mathrm{diag}(\sigma_{0},...,\sigma_{0})$. The prior is the “twisted-normal” prior of Haario et al. (1999), with density function proportional to $p(\bm{\theta})\propto\exp\biggl{\{}-\theta_{1}^{2}/200-(\theta_{2}-b\theta_{1}^{2}+100b)^{2}/2-\mathbbm{1}_{\{d_{\theta}>2\}}\sum_{j=3}^{d_{\theta}}\theta_{j}^{2}/2\biggr{\}},$ where $\mathbbm{1}_{B}$ denotes the indicator function of the set $B$. This prior is essentially a product of independent Gaussian distributions with the exception that the component for $(\theta_{1},\theta_{2})$ is modified to produce a “banana shape”, with the strength of the bivariate dependence determined by the parameter $b$. Simulation from $p(\bm{\theta})$ is achieved by first drawing $\bm{\theta}$ from a $d_{\theta}$-dimensional multivariate Gaussian as $\bm{\theta}\sim\mathcal{N}(\bm{0},\bm{A})$, where $\bm{A}=\mathrm{diag}(100,1,...,1)$, and then placing the value $\theta_{2}+b\theta_{1}^{2}-100b$ in the slot for $\theta_{2}$. We set $b=0.1$, a value inducing a strong correlation in the prior between the first two components of $\bm{\theta}$. We consider $d_{\theta}=20$, i.e. both $\bm{y}$ and $\bm{\theta}$ have length 20, and (same as Li et al., 2017; Picchini & Tamborrino, 2022) we set $\sigma_{0}=1$ as fixed and known, with observations given by the 20-dimensional vector $\bm{y}_{o}=(10,0,...,0)$, where only the first entry is non-zero.
@@
(a) Prior $p(\theta_{1},\theta_{2})$ (black) and exact posterior (green)
(b) Prior $p(\theta_{1},\theta_{2})$ (black) and draws from SeMPLE (brown)
Figure 54: Twisted prior example: prior, true posterior and draws from SeMPLE with $r=1.$

### v3

@@ G.3 Results when using $R=2$ for all algorithms
Figure 37: Bernoulli GLM model. Posterior plots obtained with SeMPLE and obtained at round $r=1$ after 5,000 model simulations, and at $r=2$ after 10,000 model simulations.
@@ G.4 Results when using $R=10$ for SNL/SNPE-C and $R=2$ for SeMPLE
We executed the experiment anew, but this time by allocating the $10^{4}$ model simulations uniformly across 2 rounds for SeMPLE and 10 rounds for SNL and SNPE-C (as per SBIBM defaults). Results are in Figure 40 and Figure 41. Once more, we recommend the reader to consider the Wasserstein distances and not C2ST for this example, since that posterior is unimodal, as motivated in the previous section. The distances show that on average SeMPLE is as accurate as SNL but not before $10^{4}$ model simulations are executed, while SNPE-C shows a large variability in the results. Once more, SeMPLE comes with the benefit of a much faster inference (Figure 41): in this case SeMPLE is 26 times faster than SNPE-C and 74 times faster than SNL.
@@ Appendix H Ornstein-Uhlenbeck
This makes the Ornstein-Uhlenbeck model one of the very few examples of SDEs where it is trivial to evaluate the exact likelihood function. The likelihood function for the parameters $\bm{\theta}=(\alpha,\beta,\sigma)$ is written as (due to the Markovianity of the solution of an SDE)
@@ H.3 Results when using $R=4$ for all algorithms
In terms of runtime, Figure 45 shows that SeMPLE has a much much smaller runtime compared to both SNL and SNPE-C, even if were to include the 16 minutes of BIC computations into account. Especially SNPE-C has much longer runtimes with this model, combined with the worst performance metric results out of all three algorithms. Plots of the posterior marginals and posterior draws from the last round (r=4) of each algorithm are in Figures 46-47, for each of the 10 independent runs. Acceptance rates of the Metropolis-Hastings step of SeMPLE are in Figure 48, showing that, particularly for the third and fourth round, the acceptance rate is fairly high (compatibly with being an MCMC algorithm).
@@ I.2 Inference setup
Prior to running the inference, for each of the nine summary statistics we computed trimmed means, where $1.25\%$ of the values were trimmed from each end and corresponding trimmed standard deviations, obtained from 10,000 prior predictive simulations, and we used these to standardise both the observed summaries and the summaries simulated during the inference. For each SeMPLE run we set a budget of 30 0003000030\,00030 000 model simulations, uniformly distributed across 3 algorithm rounds. The number of mixture components is set to $K=15$ (a number which was picked without any specific tuning) and we set the covariance matrices in GLLiM to be full unconstrained matrices, all different across the $K$ components. Moreover, we do not inflate the covariance in the Metropolis-Hastings step (ie we set $\gamma=1$). The mixture probability threshold to remove negligible mixture components during the SeMPLE run was set to $0.005$.
For a MJP Lotka-Volterra model, we did not find a ready implementation in SBIBM (there, an ODE-based Lotka-Volterra model is provided, not a MJP version), and this is likely because with such a model it is impossible to perfectly control the computational effort, as for a given value of $\theta$ the time spent to complete a trajectory is random. For this reason, we found more practical to compare SeMPLE with approximate Bayesian computation (ABC) experiments using the code from Picchini & Tamborrino (2024). For SMC-ABC, we used the same observed data and the same standardised summary statistics as with SeMPLE. We consider sequential Monte Carlo ABC (SMC-ABC), and also use the posterior from SMC-ABC as “reference posterior”. SMC-ABC is the state-of-art approach to ABC inference and is the algorithm of choice in recent inference platforms such as ABCpy (Dutta et al., 2017) and pyABC (Schälte et al., 2022). We consider two versions of SMC-ABC that differ in terms of proposal function: the first one is the “classic” SMC-ABC, in that it uses as proposal distribution a Gaussian random walk of the type $\theta^{**}\sim N(\theta^{*},2\Sigma)$ (Beaumont et al., 2009, Filippi et al., 2013), where $\theta^{*}$ is a parameter “particle” randomly sampled from the most recent set of “parameter particles” (these providing an ABC approximation to the posterior), and $\Sigma$ is the empirical weighted covariance matrix of these “parameter particles”. This first sampler is named standard in Picchini & Tamborrino (2024), as it is the most common SMC-ABC proposal sampler, as found in most literature and in available software packages. A second SMC-ABC sampler is the one denoted blockedopt in Picchini & Tamborrino (2024), which is also a Gaussian proposal sampler but not of random-walk type, and is much more resource-effective, as it more rapidly “guide” the parameters towards the bulk of the posterior region (see Picchini & Tamborrino, 2024 for details). We ran both SMC-ABC versions with 1,000 particles.
@@ I.3 Results
We now consider a model with 20 parameters and a challenging posterior characterised by a strong correlation between some of the parameters. This case study is particularly interesting, as the likelihood only provides location information about the unknown parameter while the dependence structure in the posterior comes mostly from the prior. For this reason, the posterior dependence changes direction depending on whether the likelihood locates the posterior in the left or right tail of the prior (see Nott et al., 2018 for a graphical illustration). This case study was analysed in an ABC context in Li et al. (2017) and Picchini & Tamborrino (2024). The model assumes $\bm{y}=(y_{1},...,y_{d_{\theta}})$ drawn from a $d_{\theta}$-dimensional Gaussian $\bm{y}\sim\mathcal{N}(\bm{\theta},\bm{\Psi})$, with $\bm{\theta}=(\theta_{1},...,\theta_{d_{\theta}})$ and diagonal covariance matrix $\Psi=\mathrm{diag}(\sigma_{0},...,\sigma_{0})$. The prior is the “twisted-normal” prior of Haario et al. (1999), with density function proportional to $p(\bm{\theta})\propto\exp\biggl{\{}-\theta_{1}^{2}/200-(\theta_{2}-b\theta_{1}^{2}+100b)^{2}/2-\mathbbm{1}_{\{d_{\theta}>2\}}\sum_{j=3}^{d_{\theta}}\theta_{j}^{2}/2\biggr{\}},$ where $\mathbbm{1}_{B}$ denotes the indicator function of the set $B$. This prior is essentially a product of independent Gaussian distributions with the exception that the component for $(\theta_{1},\theta_{2})$ is modified to produce a “banana shape”, with the strength of the bivariate dependence determined by the parameter $b$. Simulation from $p(\bm{\theta})$ is achieved by first drawing $\bm{\theta}$ from a $d_{\theta}$-dimensional multivariate Gaussian as $\bm{\theta}\sim\mathcal{N}(\bm{0},\bm{A})$, where $\bm{A}=\mathrm{diag}(100,1,...,1)$, and then placing the value $\theta_{2}+b\theta_{1}^{2}-100b$ in the slot for $\theta_{2}$. We set $b=0.1$, a value inducing a strong correlation in the prior between the first two components of $\bm{\theta}$. We consider $d_{\theta}=20$, i.e. both $\bm{y}$ and $\bm{\theta}$ have length 20, and (same as Li et al., 2017; Picchini & Tamborrino, 2024) we set $\sigma_{0}=1$ as fixed and known, with observations given by the 20-dimensional vector $\bm{y}_{o}=(10,0,...,0)$, where only the first entry is non-zero.
@@
(a) Draws from prior $p(\theta_{1},\theta_{2})$ (black) and exact posterior (green)
(b) Draws from prior $p(\theta_{1},\theta_{2})$ (black) and from SeMPLE (brown)
Figure 55: Twisted prior example: prior, true posterior and draws from SeMPLE at round $r=1.$
@@
Appendix K Biological model of the translation kinetics after mRNA transfection
This model serves as an example of a multidimensional SDE model with a higher number of model parameters, compared to the Ornstein-Uhlenbeck model. Moreover here the model is challenging as it is partially observed (only one coordinate is observed), and observations incorporate measurement measurement error. Reference posterior samples cannot be obtained exactly, as the transition densities are not known for this model.
K.1 Model definition
We consider the SDE model from Pieschner et al. (2022) in a simulation study. In Pieschner et al. (2022) the model was used to describe translation kinetics after mRNA transfection. Specifically, in a real data scenario, time lapse microscopy images of the cells are taken for several hours, and for the first hour, the cells would be incubated with mRNA lipoplexes. Released mRNA gets translated into a green fluorescent protein (GFP) which causes the cells to fluoresce. The production of GFPs is the translation process. The SDE below is used to study the translation kinetics of one cell, based on one observed fluorescence trajectory, where the two-dimensional stochastic process $(m(t),p(t))$ has $m(t)$ to represent the amount of mRNA molecules at time $t$, and $p(t)$ is the amount of GFP molecules at time $t$. The SDE model is given by
$d\begin{pmatrix}m\\ p\end{pmatrix}(t)=\begin{pmatrix}-\delta\cdot m(t)\\ k\cdot m(t)-\gamma\cdot p(t)\end{pmatrix}dt+\begin{pmatrix}\sqrt{\delta\cdot m(t)}&0\\ 0&\sqrt{k\cdot m(t)+\gamma\cdot p(t)}\end{pmatrix}dB_{t},$ (35)
where $dB_{t}$ is a two-dimensional standard Brownian motion. It is assumed that all mRNA molecules (within one cell) are released at once from the lipoplexes and denote this initial time point by $t_{0}$. Before $t_{0}$, there are neither mRNA nor GFP molecules, and at $t_{0}$, an amount $m_{0}$ of mRNA molecules is released, i.e
$\begin{pmatrix}m(t)\\ p(t)\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix}\qquad\text{for $t<t_{0}$}$
while $m(t_{0})=m_{0}$ and $p(0)=0$.
We take as observable mapping
${y}(t_{i})=\log(\textrm{scale}\cdot p(t_{i})+\textrm{offset})+\varepsilon(t_{i})\qquad i=1,...,n,$
where $\varepsilon(t_{i})$ is iid Gaussian measurement error $\varepsilon(t_{i})\sim\mathcal{N}(0,\sigma^{2})$, and $t_{1},...,t_{n}$ are observational time instants (see further below). Note that the observations depend only indirectly on process $\{m(t)\}$, and only $\{p(t)\}$ is observed. The parameter vector we wish to infer is $\bm{\theta}=(\delta,\gamma,k,m_{0},\textrm{scale},t_{0},\textrm{offset},\sigma)$.
It has been shown by Pieschner et al. (2022) that SDE modelling improves identifiability of parameters, compared to using a corresponding ODE model. Nevertheless, some parameters are still unidentifiable, see in particular supplementary section A.4.1 in Pieschner et al. (2022).
Solutions to the SDE (35) are simulated using an Euler-Maruyama scheme implemented in Rcpp (Eddelbuettel et al., 2024) from $t=t_{0}$ to $t=30$, with step size 0.01. The observed time series is interpolated from the Euler-Maruyama approximation at $n=60$ equidistant time points from $t_{1}=0.5$ to $t_{n}=30$. Note that this implies that observations can be made prior to $t=t_{0}$ where $p(t_{i})=0$, $t_{i}<t_{0}$.
Inference is performed on the parameters log-scale, and the prior distributions are all Gaussian as follows
$\log\delta$ $\sim\mathcal{N}(-0.694,0.5),\qquad\log\gamma\sim\mathcal{N}(-3,0.5)$
$\log k$ $\sim\mathcal{N}(0.027,0.5),\qquad\log m_{0}\sim\mathcal{N}(5.704,0.5)$
$\log\textrm{scale}$ $\sim\mathcal{N}(0.751,0.5),\qquad\log t_{0}\sim\mathcal{N}(-0.164,0.5),$
$\log\textrm{offset}$ $\sim\mathcal{N}(2.079,0.5),\qquad\log\sigma\sim\mathcal{N}(-2,0.5).$
SeMPLE, SNL and SNPE-C were run for five different observed data sets, each generated with a different set of parameters given in Table 3. For each method, the total simulation budget was set to 30k model simulations that were uniformly distributed across the number of algorithm rounds. For SeMPLE, the number of algorithm rounds was set to $R=3$, the number of starting mixture components was set to $K=10$, and the covariance matrices in GLLiM were set to be full unconstrained matrices. We did not inflate the covariance in the Metropolis-Hastings step ($\gamma=1$). The mixture probability threshold to remove negligible mixture components was set to 0.005. For SNPE-C and SNL the number of algorithm rounds was set to 10.
$\log\delta$ $\log\gamma$ $\log k$ $\log m_{0}$ $\log\textrm{scale}$ $\log t_{0}$ $\log\textrm{offset}$ $\log\sigma$
Data set 1 -0.694 -3 0.027 5.704 0.751 -0.164 2.079 -2
Data set 2 -1.827 -3.138 -0.426 6.869 0.697 0.465 2.787 -1.993
Data set 3 -1.153 -3.501 -0.181 4.601 0.386 -2.121 2.185 -2.578
Data set 4 -1.297 -2.134 1.118 5.577 0.938 0.12 1.122 -2.854
Data set 5 -0.243 -4.596 0.168 5.573 -0.086 1.234 1.392 -1.655
Table 3: Translation kinetics model: parameter values used to generate observations for the five data sets.
K.2 Results
Figures 57-61 compare the marginal posteriors obtained by SeMPLE, SNPE-C and SNL, where each figure corresponds to a different dataset. All results are obtained after 30k model simulations but the number of algorithm rounds is 10 for both SNPE-C and SNL, while we ran 3 rounds for SeMPLE. Inference is very similar across the three methods, however it is clear that only SeMPLE can identify $t_{0}$ accurately, while SNL and SNPE-C provide a biased estimation of $t_{0}$ (an exception being the inference for dataset 3 in Figure 59, where we deliberately challenged the inference by setting the ground truth value for $t_{0}$ far in the left tail of the prior). Moreover, SNPE-C is often unable to infer the measurement error variability $\sigma$ (datasets 2, 4, 5). As previously found in Pieschner et al. (2022), the lack of identifiability of parameters $(k,m_{0},\textrm{scale})$ is confirmed in our analyses. However the other parameters are correctly inferred by SeMPLE. Finally, in Figures 62-66 we focus on displaying the performance of SeMPLE for each round.
Figure 57: Translation kinetics model, observed data set #1 out of 5: marginal posteriors from SeMPLE, SNPE-C and SNL (after 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines). For SNPE-C and SNL we show results after 10 algorithm rounds, and after 3 rounds for SeMPLE.
Figure 58: Translation kinetics model, observed data set #2 out of 5: marginal posteriors from SeMPLE, SNPE-C and SNL (after 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines). For SNPE-C and SNL we show results after 10 algorithm rounds, and after 3 rounds for SeMPLE.
Figure 59: Translation kinetics model, observed data set #3 out of 5: marginal posteriors from SeMPLE, SNPE-C and SNL (after 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines). For SNPE-C and SNL we show results after 10 algorithm rounds, and after 3 rounds for SeMPLE.
Figure 60: Translation kinetics model, observed data set #4 out of 5: marginal posteriors from SeMPLE, SNPE-C and SNL (after 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines). For SNPE-C and SNL we show results after 10 algorithm rounds, and after 3 rounds for SeMPLE.
Figure 61: Translation kinetics model, observed data set #5 out of 5: marginal posteriors from SeMPLE, SNPE-C and SNL (after 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines). For SNPE-C and SNL we show results after 10 algorithm rounds, and after 3 rounds for SeMPLE.
Figure 62: Translation kinetics model, observed data set #1 out of 5: marginal posteriors from SeMPLE for every algorithm round (corresponding to 10k, 20k and 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines).
Figure 63: Translation kinetics model, observed data set #2 out of 5: marginal posteriors from SeMPLE for every algorithm round (corresponding to 10k, 20k and 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines).
Figure 64: Translation kinetics model, observed data set #3 out of 5: marginal posteriors from SeMPLE for every algorithm round (corresponding to 10k, 20k and 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines).
Figure 65: Translation kinetics model, observed data set #4 out of 5: marginal posteriors from SeMPLE for every algorithm round (corresponding to 10k, 20k and 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines).
Figure 66: Translation kinetics model, observed data set #5 out of 5: marginal posteriors from SeMPLE for every algorithm round (corresponding to 10k, 20k and 30k model simulations), priors (black lines) and ground truth parameters (vertical red lines).